# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Evasion Adversarial Attacks on ML-NIDS Author’s draft for soliciting feedback: 5 Nov. 2024

### v2

@@ Abstract
As network security threats continue to evolve, safeguarding Machine Learning (ML)-based Network Intrusion Detection Systems (NIDS) from adversarial attacks is crucial. This paper introduces the notion of feature perturb-ability and presents a novel Perturb-ability Score (PS) metric that identifies NIDS features susceptible to manipulation in the problem-space by an attacker. By quantifying a feature’s susceptibility to perturbations within the problem-space, the PS facilitates the selection of features that are inherently more robust against evasion adversarial attacks on ML-NIDS during the feature selection phase. These features exhibit natural resilience to perturbations, as they are heavily constrained by the problem-space limitations and correlations of the NIDS domain. Furthermore, manipulating these features may either disrupt the malicious function of evasion adversarial attacks on NIDS or render the network traffic invalid for processing (or both).
This proposed novel approach employs a fresh angle by leveraging network domain constraints as a defense mechanism against problem-space evasion adversarial attacks targeting ML-NIDS.
We demonstrate the effectiveness of our PS-guided feature selection defense in enhancing NIDS robustness. Experimental results across various ML-based NIDS models and public datasets show that selecting only robust features (low-PS features) can maintain solid detection performance while significantly reducing vulnerability to evasion adversarial attacks. Additionally, our findings verify that the PS effectively identifies NIDS features highly vulnerable to problem-space perturbations.
@@ 1 Introduction
In the contemporary landscape of cybersecurity, Machine Learning (ML) has emerged as a cornerstone technology for Network Intrusion Detection Systems (NIDS). The ability of ML algorithms to process and classify vast amounts of network traffic data with high accuracy makes them indispensable for identifying malicious activities and safeguarding network integrity [1] [2]. However, the effectiveness of ML-based systems is challenged by adversarial attacks, which exploit vulnerabilities in the ML models to fool their detection mechanisms [3]. There are multiple types of adversarial attacks, such as poisoning [4], backdoor [5], model stealing [6], and evasion attacks, which are the main focus of this paper.
@@ 1.1 Feature-Space vs Problem-Space Evasion Adversarial Attacks Against ML-NIDS
It is crucial to differentiate between feature-space and problem-space (real-world objects) in ML [7]. Feature-space refers to the representation of input data as features used by the ML model, while problem-space refers to the raw input data before feature extraction. Ibitoye et al. [3] introduced the concept of ’space’ in their taxonomy of adversarial attacks for network security, distinguishing between feature-space and problem-space attacks. Feature-space adversarial attacks manipulate or perturb feature vectors, assuming that the attacker can directly access and alter these features; however, this assumption is often unrealistic in practical scenarios. In contrast ,problem-space adversarial attacks modify or perturb actual data, such as network traffic (e.g., adding delays) or packets (e.g., padding the payload), which attackers typically can access. Fig.  1 illustrates this distinction.
@@ 1.2 Perturb-ability of Features in Problem-Space Against NIDS
[High Perturb-ability, L2 [Features can be perturbed in problem-space while maintaining NIDS constrains, L2]] [Medium Perturb-ability, L3[Features with medium possibility to be perturbed in problem-space while maintaining NIDS constrains,L3]] [Low Perturb-ability, L4[Features extremely hard to perturb in problem-space while maintaining NIDS constrains,L4]] ]
@@ 1.4 Contributions
2 Threat Model of a “Practical” Evasion Adversarial Attack Against NIDS
In this section, we describe what we consider a threat model for a practical evasion adversarial attack against NIDS.
Despite the critical importance of defending against adversarial attacks, much of the existing research has focused on computer vision datasets, where the primary objective is to fool human perception [20]. This focus has led to the development of algorithms optimized for visual data, which may not directly translate to the network security domain. Furthermore, many studies on adversarial attacks against ML-NIDS operate under the assumption that adversaries have full control over the feature-space, an unrealistic scenario in practical settings [7].
Attacker’s Knowledge: In a practical setting, we assume that the attacker has no knowledge (black-box attack) of the following: training data, selected features (feature vector), feature extractor, feature transformers, ML algorithm, activation functions, parameters, hyperparameters, and confidence intervals. Moreover, we assume that the attackers have no access to model outputs (i.e., they cannot query the model).
Attacker’s Capability: We assume that the attacker has no access to the feature vector. In this scenario, the attacker has access to and can only alter network packets (problem-space attack). To bypass detection, they must craft these adversarial packets while adhering to several constraints: (1) Maintaining Functionality: Malicious capability must be preserved alongside the network functionality of the packets. (2) NIDS Feature Constraints: The attacker needs to consider the feature limitations of the NIDS model. These constraints might involve specific data formats, packet sizes, correlations, or limitations on certain network protocols.
Attacker’s Goal: The attacker’s main objective is to compromise the integrity of the NIDS by evading detection (evasion adversarial attacks) while maintaining malicious and network functionalities of the perturb-ed network flow.
Such a threat model might be infeasible, as recent research studies have cast doubt on the practicality of evasion adversarial attacks against NIDS [1] [21] [22]. However, in this paper, we focus solely on the possibility and implications of adding perturbations to NIDS features in the problem-space. Therefore, we will assume that such a threat model is feasible for an extremely powerful attacker.
@@ 3.2 Evaluating PS
Figure 6: Visualization of the PS3[$f_{i}$] Evaluation Equation
@@ 3.2.2 PS2[$f_{i}$]: The range of Possible Values of a Feature
PS2[$f_{i}$] will be 1 if $f_{i}$’s number of Possible Values ($PV$) is greater than 255 (this feature will be similar to computer vision’s pixel, and it will be flexible to perturb). On the other hand, if $f_{i}$’s $PV$ ($PV$[$f_{i}$]) is less than or equal to 255, PS2[$f_{i}$] will be equal to a linear function where its output is 1 if $f_{i}$’s $PV$ is 255, and 0.5 if $f_{i}$’s $PV$ is 2 (binary). If $PV[f_{i}]$ is less than 2 (equals 1), it indicates that $f_{i}$ is non-perturb-able, in which case $\text{PS}_{2}[f_{i}]$ will be set to 0. However, in this case (where $PV[f_{i}]$ equals 1), we recommend dropping that feature, as it does not contribute meaningful information to the ML model. Fig. 5 shows the visualization of the PS2[$f_{i}$] evaluation equation.
@@ 3.2.3 PS3[$f_{i}$]: Correlated Features
PS3[$f_{i}$] will follow a linear function, where its output is 0.5 if the number of Correlated Features ($CF$) of $f_{i}$ is equal to or greater than a threshold (the maximum number observed in our experiments was 10, which we chose as the threshold), and 1 if $f_{i}$’s $CF$ (CF[$f_{i}$]) is 0. Fig. 6 shows the visualization of the PS3[$f_{i}$] evaluation equation.
@@ 3.2.4 PS4[$f_{i}$]: Features that attackers cannot access
3.2.5 PS5[$f_{i}$]: Features Correlated with numerous flow Packets
This PS field considers features that are correlated with numerous flow packets.
@@ 4.1 Datasets
For our experiments, we used 2 public NIDS datasets;
(1) The UNSW-NB15 dataset [25] is a comprehensive network intrusion detection dataset designed to address the limitations of older benchmarks, such as KDD99 and NSL-KDD [26]. It contains 2,540,044 records, divided into 2,218,761 normal and 321,283 attack samples, spanning 9 different attack categories. The dataset includes 49 features: 47 describing the flow and 2 label features (attack_cat and Label). The UNSW-NB15 features capture various aspects of network behavior, including basic, content, time, and connection-based attributes. It was generated using the IXIA PerfectStorm tool in a controlled testbed environment, simulating both normal network activities and contemporary attack scenarios. This approach ensures that the dataset reflects modern network traffic patterns and attack vectors, making it a valuable resource for evaluating and developing network intrusion detection systems in today’s cybersecurity landscape.
(2) The CSE-CIC-IDS2018 dataset [27] is a significant resource for network security research, capturing realistic network traffic over five days and simulating various attack scenarios, including Brute Force, DoS, Botnets, Infiltration, and DDoS, interspersed with benign traffic. With over 80 features extracted from the captured traffic using CICFlowMeter, providing detailed packet and flow-level information, it is widely applicable to machine learning tasks such as classification and anomaly detection. CSE-CIC-IDS2018 is currently one of the most up-to-date datasets and has been shown to improve ML model accuracy in both training and testing [2]. However, recent studies [28] have identified several significant issues related to feature generation and labeling in the CSE-CIC-IDS2018 dataset. Therefore, we utilized the improved version of CSE-CIC-IDS2018 created by Liu et al. [28], which addresses these issues by updating the dataset, modifying the CICFlowMeter Tool, and adding new features like total TCP flow duration, Fwd/Bwd RST flag counts, and support for ICMP protocol.
@@ 4.2 Features’ Analysis and Correlations
TABLE I: The number and percentage of features in every perturb-ability class, based on our proposed PS, where green indicates low perturb-ability features class, yellow indicates medium perturb-ability features class, and red indicates high perturb-ability features class
@@
Figure 11: The histogram of PS values for each dataset where green indicates low perturb-ability features class, yellow indicates medium perturb-ability features class, and red indicates high perturb-ability features class
TABLE II: UNSW-NB15 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
srcip sport dstip dsport proto state dur
sbytes dbytes sttl dttl sloss dloss service
Sload Dload Spkts Dpkts swin dwin stcpb
dtcpb smeansz dmeansz trans_depth res_bdy_len Sjit Djit
Stime Ltime Sintpkt Dintpkt tcprtt synack ackdat
is_sm_ips_ports ct_state_ttl ct_flw_http_mthd is_ftp_login ct_ftp_cmd ct_srv_src ct_srv_dst
ct_dst_ltm ct_src_ ltm ct_src_dport_ltm ct_dst_sport_ltm ct_dst_src_ltm
TABLE III: CSE-CIC-IDS2018 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
Flow ID Src IP Src Port Dst IP Dst Port
Protocol Timestamp Flow Duration Total Fwd Packet Total Bwd packets
Total Len of Fwd Pack Total Len of Bwd Pack Fwd Packet Length Max Fwd Packet Length Min Fwd Packet Length Mean
Fwd Packet Length Std Bwd Packet Length Max Bwd Packet Length Min Bwd Packet Length Mean Bwd Packet Length Std
Flow Bytes/s Flow Packets/s Flow IAT Mean Flow IAT Std Flow IAT Max
Flow IAT Min Fwd IAT Total Fwd IAT Mean Fwd IAT Std Fwd IAT Max
Fwd IAT Min Bwd IAT Total Bwd IAT Mean Bwd IAT Std Bwd IAT Max
Bwd IAT Min Fwd PSH Flags Bwd PSH Flags Fwd URG Flags Bwd URG Flags
Fwd RST Flags Bwd RST Flags Fwd Header Length Bwd Header Length Fwd Packets/s
Bwd Packets/s Packet Length Min Packet Length Max Packet Length Mean Packet Length Std
Packet Len Variance FIN Flag Count SYN Flag Count RST Flag Count PSH Flag Count
ACK Flag Count URG Flag Count CWR Flag Count ECE Flag Count Down/Up Ratio
Average Packet Size Fwd Segment Size Avg Bwd Segment Size Avg Fwd Bytes/Bulk Avg Fwd Packet/Bulk Avg
Fwd Bulk Rate Avg Bwd Bytes/Bulk Avg Bwd Packet/Bulk Avg Bwd Bulk Rate Avg Subflow Fwd Packets
Subflow Fwd Bytes Subflow Bwd Packets Subflow Bwd Bytes FWD Init Win Bytes Bwd Init Win Bytes
Fwd Act Data Pkts Fwd Seg Size Min Active Mean Active Std Active Max
Active Min Idle Mean Idle Std Idle Max Idle Min
ICMP Code ICMP Type Total TCP Flow Time
4.3 Pre-processing
@@ 4.4 Machine and Deep Learning Models
(1) The Vanilla Neural Network Model comprises three hidden layers with 64, 32, and 16 neurons, respectively, each employing the ReLU activation function to capture non-linear relationships. The output layer features a single neuron with a sigmoid activation function, suited for binary classification tasks. The model is optimized using the Adam optimizer, minimizing binary cross-entropy loss. The network is trained over 10 epochs with a batch size of 32.
(2) The SVM model employs a linear kernel for binary classification. It is trained on the resampled dataset to identify optimal decision boundaries between classes. The linear SVM aims to maximize the margin between support vectors while minimizing classification errors.
(3) The Random Forest model consists of an ensemble of 100 decision trees, designed for binary classification tasks. Each tree is trained on a random subset of the data, enhancing the model’s robustness and reducing overfitting. The aggregation of predictions from individual trees leads to improved accuracy and stability in classification outcomes. The Random Forest model was also used for feature importance analysis. The model quantifies the contribution of each feature to the model’s predictive performance. A DataFrame is created to rank the features and their corresponding importance scores.
(4) The CNN model is designed for sequential data classification. The input data is reshaped to accommodate the CNN architecture, with a shape of (samples, timesteps, features). The model consists of a 1D convolutional layer with 64 filters and a kernel size of 3, followed by a max-pooling layer to reduce dimensionality. The flattened output is then passed through a dense layer with 100 neurons and ReLU activation, culminating in an output layer with a softmax activation function to produce class probabilities. The model is trained over 10 epochs with a batch size of 32.

### v3

# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Problem-Space Evasion Adversarial Attacks on Flow-based ML-NIDS

@@ Abstract
As network security threats continue to evolve, safeguarding flow-based Machine Learning (ML)-based Network Intrusion Detection Systems (NIDS) from adversarial attacks is crucial. This paper introduces the notion of feature perturb-ability and presents a novel Perturb-ability Score (PS) metric, which identifies NIDS features that are susceptible to manipulation in the problem-space by an attacker. By quantifying a feature’s susceptibility to perturbations within the problem-space, the PS facilitates the selection of features that are inherently more robust against evasion adversarial attacks on flow-based ML-NIDS during the feature selection phase. These features exhibit natural resilience to perturbations as they are heavily constrained by the problem-space limitations and correlations specific to the NIDS domain. Furthermore, manipulating these features may either disrupt the malicious functionality of evasion adversarial attacks on NIDS or render the network traffic invalid for processing (or both). This novel approach introduces a fresh perspective by leveraging problem-space NIDS domain constraints as a defense mechanism against problem-space evasion adversarial attacks targeting flow-based ML-NIDS.
We demonstrate the effectiveness of our PS-guided feature selection defense in enhancing flow-based NIDS robustness. Experimental results across various ML-based NIDS models and public datasets show that selecting only robust features (low-PS features) can maintain solid detection performance while significantly reducing vulnerability to evasion adversarial attacks. Additionally, our findings verify that the PS effectively identifies flow-based NIDS features highly vulnerable to problem-space perturbations.
@@ 1 Introduction
In the modern cybersecurity landscape, flow-based Machine Learning (ML) has emerged as a cornerstone technology for Network Intrusion Detection Systems (NIDS). The ability of ML algorithms to process and classify traffic at the flow level, where a flow represents aggregated packet metadata without inspecting payload content, with high speed and accuracy makes them indispensable for identifying malicious activities and safeguarding network integrity [35] [5]. However, the effectiveness of ML-based systems is challenged by adversarial attacks, which exploit vulnerabilities in the ML models to fool their detection mechanisms [25]. There are multiple types of adversarial attacks, such as poisoning [15], backdoor [32], model stealing [26], and evasion attacks, the latter being the primary focus of this paper.
@@ 1.1 Feature-Space vs Problem-Space Evasion Adversarial Attacks Against ML-NIDS
It is crucial to differentiate between feature-space and problem-space (real-world objects) in ML [10]. Feature-space refers to the representation of input data as features used by the ML model, while problem-space refers to the raw input data before feature extraction. Ibitoye et al. [25] introduced the concept of ’space’ in their taxonomy of adversarial attacks for network security, distinguishing between feature-space and problem-space attacks. Feature-space adversarial attacks manipulate or perturb feature vectors, assuming that the attacker can directly access and alter these features; however, this assumption is often unrealistic in practical scenarios. In contrast, problem-space adversarial attacks modify or perturb actual raw input data, such as network traffic flows (e.g., adding delays) or packets (e.g., padding the payload), which attackers can typically access. Fig. 1 illustrates this distinction.
@@ 1.2 Perturb-ability of Features in Problem-Space Against NIDS
[High
Perturb-ability, L2 [Features can be perturbed in problem-space
while maintaining NIDS constrains, L2]] [Medium
Perturb-ability, L3[Features with medium possibility to be perturbed in
problem-space while maintaining NIDS constrains,L3]] [Low
Perturb-ability, L4[Features extremely hard to perturb in problem-space
while maintaining NIDS constrains,L4]] ]
@@ 1.4 Contributions
2 Threat Model
The following is our threat model; Target: We assume that the target model is a flow-based NIDS, rather than packet-based or raw traffic-based NIDS. Attacker’s Knowledge: We assume that the attacker has no knowledge of the selected features (feature vector) used by the model. However, other information, such as the ML algorithm and activation functions, may be known to the attacker. Attacker’s Capability: We assume that the attacker does not have access to the feature vector but can only access and alter network packets or flows (problem-space modifications). In the NIDS context, perturbing network flow at a packet granular level may be challenging if it results in a network flow that does not correspond to any valid network behavior. Thus, we assume that attackers might deploy innovative approaches to crafting problem-space attacks by changing flow behavior or manipulating hosts’ behavior instead of directly modifying packets, for example, adjusting the output rate, adding packet delay, introducing packet loss, and implementing duplication [12]. Additionally, attackers must be capable of creating adversarial flows while adhering to several constraints: (1) Maintaining Functionality: The malicious capability must be preserved alongside the network functionality of the packets. (2) NIDS Feature Constraints: The attacker needs to consider the feature limitations of the NIDS model. These constraints might involve specific data formats, packet sizes, correlations, or limitations on certain network protocols. Attacker’s Goal: The attacker aims to compromise the integrity of the NIDS by evading detection (evasion adversarial attacks) while maintaining the malicious functionality and operational integrity of the perturbed network flow.
Although recent research has cast doubt on the practicality of evasion adversarial attacks against NIDS [35, 7, 16], since knowing the ML algorithm, activation functions, or querying the NIDS is unlikely, this paper nonetheless focuses on exploring the possibility and implications of adding perturbations to NIDS features in the problem space.
@@ 3.2 Evaluating PS
(a) Visualization of the PS2[$f_{i}$] Evaluation Equation
(b) Visualization of the PS3[$f_{i}$] Evaluation Equation
Figure 5: Visualizations of the PS2[$f_{i}$] and PS3[$f_{i}$] Evaluation Equations
@@ 3.2.2 PS2[$f_{i}$]: The range of Possible Values of a Feature
PS2[$f_{i}$] will be 1 if $f_{i}$’s number of Possible Values ($PV$) is greater than 255 (this feature will be similar to computer vision’s pixel, and it will be flexible to perturb). On the other hand, if $f_{i}$’s $PV$ ($PV$[$f_{i}$]) is less than or equal to 255, PS2[$f_{i}$] will be equal to a linear function where its output is 1 if $f_{i}$’s $PV$ is 255, and 0.5 if $f_{i}$’s $PV$ is 2 (binary). If $PV[f_{i}]$ is less than 2 (equals 1), it indicates that $f_{i}$ is non-perturb-able, in which case $\text{PS}_{2}[f_{i}]$ will be set to 0. However, in this case (where $PV[f_{i}]$ equals 1), we recommend dropping that feature, as it does not contribute meaningful information to the ML model. Fig. 5a shows the visualization of the PS2[$f_{i}$] evaluation equation.
@@ 3.2.3 PS3[$f_{i}$]: Correlated Features
PS3[$f_{i}$] will follow a linear function, where its output is 0.5 if the number of Correlated Features ($CF$) of $f_{i}$ is equal to or greater than a threshold (the maximum number observed in our experiments was 10, which we chose as the threshold), and 1 if $f_{i}$’s $CF$ (CF[$f_{i}$]) is 0. Fig. 5b shows the visualization of the PS3[$f_{i}$] evaluation equation.
@@ 3.2.4 PS4[$f_{i}$]: Features that attackers cannot access
3.2.5 PS5[$f_{i}$]: Flow-wide Correlation
This PS field considers whether modifying a feature requires altering the entire flow. Features like mean or standard deviation, which depend on multiple packets within a flow, are harder to perturb due to their broader impact on the network flow.
@@ 4.1 Datasets
For our experiments, we used 2 public NIDS datasets; (1) UNSW-NB15 [29]: A widely-used NIDS dataset with 2.5M records (2.2M normal, 0.32M attacks) across 9 attack categories. It includes 49 features and represents realistic network traffic and diverse attack vectors. (2) CSE-CIC-IDS2018 (Improved) [28]: A dataset with various attacks (e.g., Brute Force, DoS, DDoS) and benign traffic. We use the improved version by Liu et al. [28], which resolves issues in the original dataset [33] related to feature generation and labeling, ensuring better reliability for research.
@@ 4.2 Features’ Analysis and Correlations
4.3 Pre-processing
Some researchers, such as Arp et al. [9], caution that using the IP address as a feature might lead to spurious correlations or false associations, as the model could learn to identify specific IP ranges instead of recognizing generic attack patterns. However, in our tests, we do not use the IP address as a numerical value. Instead, we extract meaningful geolocation information (e.g., country, region) to provide contextually relevant insights. This approach aims to improve real-world intrusion detection while mitigating the risks associated with direct IP-based correlations.
Table 1: The number and percentage of features in every perturb-ability class, based on our proposed PS, where green indicates low perturb-ability features class, yellow indicates medium perturb-ability features class, and red indicates high perturb-ability features class
@@ 4.3 Pre-processing
Figure 10: The histogram of PS values for each dataset where green indicates low perturb-ability features class, yellow indicates medium perturb-ability features class, and red indicates high perturb-ability features class.
@@ 4.4 Machine and Deep Learning Models
Table 2: UNSW-NB15 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
srcip sport dstip dsport proto state dur sbytes dbytes sttl dttl
sloss dloss service Sload Dload Spkts Dpkts swin dwin stcpb dtcpb
smeansz dmeansz trans_depth res_bdy_len Sjit Djit Stime Ltime Sintpkt Dintpkt tcprtt
synack ackdat is_sm_ips_ports ct_state_ttl ct_flw_http_mthd is_ftp_login ct_ftp_cmd ct_srv_src ct_srv_dst ct_dst_ltm ct_src_ ltm
ct_src_dport_ltm ct_dst_sport_ltm ct_dst_src_ltm
Table 3: CSE-CIC-IDS2018 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
Flow ID Src IP Src Port Dst IP Dst Port Protocol Timestamp
Flow Duration Total Fwd Packet Total Bwd packets Total Len of Fwd Pack Total Len of Bwd Pack Fwd Packet Length Max Fwd Packet Length Min
Fwd Packet Length Mean Fwd Packet Length Std Bwd Packet Length Max Bwd Packet Length Min Bwd Packet Length Mean Bwd Packet Length Std Flow Bytes/s
Flow Packets/s Flow IAT Mean Flow IAT Std Flow IAT Max Flow IAT Min Fwd IAT Total Fwd IAT Mean
Fwd IAT Std Fwd IAT Max Fwd IAT Min Bwd IAT Total Bwd IAT Mean Bwd IAT Std Bwd IAT Max
Bwd IAT Min Fwd PSH Flags Bwd PSH Flags Fwd URG Flags Bwd URG Flags Fwd RST Flags Bwd RST Flags
Fwd Header Length Bwd Header Length Fwd Packets/s Bwd Packets/s Packet Length Min Packet Length Max Packet Length Mean
Packet Length Std Packet Len Variance FIN Flag Count SYN Flag Count RST Flag Count PSH Flag Count ACK Flag Count
URG Flag Count CWR Flag Count ECE Flag Count Down/Up Ratio Average Packet Size Fwd Segment Size Avg Bwd Segment Size Avg
Fwd Bytes/Bulk Avg Fwd Packet/Bulk Avg Fwd Bulk Rate Avg Bwd Bytes/Bulk Avg Bwd Packet/Bulk Avg Bwd Bulk Rate Avg Subflow Fwd Packets
Subflow Fwd Bytes Subflow Bwd Packets Subflow Bwd Bytes FWD Init Win Bytes Bwd Init Win Bytes Fwd Act Data Pkts Fwd Seg Size Min
Active Mean Active Std Active Max Active Min Idle Mean Idle Std Idle Max
Idle Min ICMP Code ICMP Type Total TCP Flow Time
(1) Vanilla Neural Network: A 3-layer model (64, 32, 16 neurons) with ReLU activation, a sigmoid output layer, optimized with Adam, trained for 10 epochs (batch size 32). (2) SVM: A linear kernel SVM trained on the resampled dataset to maximize the margin between support vectors and minimize classification errors. (3) Random Forest: An ensemble of 100 trees for binary classification, reducing overfitting and providing feature importance analysis. (4) CNN: A 1D CNN with 64 filters, max-pooling, and a dense layer (100 neurons), ending with a softmax output. Trained for 10 epochs (batch size 32).